3) The equation for the instantaneous voltage across a discharging capacitor is given by the equation

๐‘ฆ = 12๐‘’ ^-x/2

a) Differentiate the function to find an equation for the gradient and calculate two gradients at x = 2 and x = 4.


4)

๐‘ฆ = 10 ๐‘™๐‘œ๐‘”3x

a)Differentiate the function to find an equation for the gradient and calculate two gradients at x = 2 and x = 8.


1
Expert's answer
2021-11-04T19:24:09-0400

3)

"\frac{dy}{dx}=12\frac{d}{dx}e^{-\frac{x}{2}}\\=12\frac{d}{dx}e^{-\frac{x}{2}}\times \frac{d}{dx}(\frac{-x}{2})" (chain rule to differentiate on)


"=12\times e^{\frac{-x}{2}}\times \frac{-1}{2}\\\frac{dy}{dx}=-6^{\frac{-x}{2}}" (equation of the radiant)


at x=2,"\frac{dy}{dx}=-6e^{\frac{-2}{2}}=\frac{-6}{e}\\"

at x=4,"\frac{dy}{dx}=-6e^{\frac{-4}{2}}=\frac{-6}{e^2}\\"

gradient at x=2 is "\frac{-6}{e}\\"

gradient at x=4 is "\frac{-6}{e^2}\\"

4)

"\frac{dy}{dx}=10\frac{d}{dx}log(3x)\\=10\times\frac{1}{ln(10)}\times \frac{1}{3x}\times 3\times 1\\\frac{dy}{dx}=\frac{10}{xln(10)}" (chain rule to differentiate on)


"\frac{dy}{dx}=\frac{10}{xln(10)}" (equation of the radiant)


at x=2,"\frac{dy}{dx}=\frac{10}{2ln(10)}=\frac{5}{ln(10)}"

at x=8,"\frac{dy}{dx}=\frac{10}{8ln(10)}=\frac{5}{4ln(10)}"

gradient at x=2 is "\frac{5}{ln(10)}\\"

gradient at x=8 is "\frac{5}{4ln(10)}\\"